# Measurement incompatibility in Bayesian multiparameter quantum estimation

### v1

@@ II.3 Optimal quantum strategies
which differs from the a priori MSL for nonvanishing prior mean, i.e., when $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,\theta_{i}\neq 0$ for all $i$; see Eq. (37) and App. A. The element $L^{ij}$ denote the $ij$-th component of $L$, and we employ the Einstein summation convention over indices $i,j=1,\dots,d$.
@@
where to obtain the rightmost expressions we have combined the estimators and the measurement by using the relabelled POVM elements [demkowicz2020multi]
@@ II.4 Lower bounds from optimal single-parameter strategies
While exact computation of $\mathcal{L}_{\mathrm{min}}$ is generally difficult, it admits a lower bound [rubio2020bayesian, Sidhu2019, demkowicz2020multi]
@@
Here and in the rest of this work, we assume that this inverse exists, which is guaranteed for a full-rank $\rho_{0}$. When this is not the case, one simply needs to to work in the relevant subspace spanned by the parametric family $\rho(\boldsymbol{\theta})$, which corresponds to the support of $\rho_{0}$. [^3] This is possible because the support the operators $\{\bar{\rho}_{j}\}$ is contained in the support of $\rho_{0}$, since $\braket{\psi|\rho_{0}|\psi}=0$ implies $\braket{\psi|\rho(\boldsymbol{\theta})|\psi}=0$ almost everywhere (in the measure-theoretic sense), which in turn implies $\braket{\psi|\bar{\rho}_{j}|\psi}=0$. This is markedly different from local QET, where the partial derivatives operators may have components also in the kernel of the state, and properties of rank-deficient models are non-trivial. For example, necessary and sufficient conditions for attainability of the SLD CRB are not fully characterized [Yang2018b, horodeckiFiveOpenProblems2022, Conlon2024c, Nurdin2024a] (except for pure states [Matsumoto2002, Pezze2017])
@@
Formally, the PM estimator is the component-wise ratio of the first-moment probability vector $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,p(x|\boldsymbol{\theta})\,\theta_{i}$ to the average probability vector $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,p(x|\boldsymbol{\theta})$. Intuitively, if $p(x|\boldsymbol{\theta})$ is formally replaced by a density operator $\rho(\boldsymbol{\theta})$, there are infinitely many inequivalent ways to “divide” $\bar{\rho}_{i}$ by $\rho_{0}$. The SPMs correspond to the symmetric choice, where “multiplication by $\rho_{0}$” is implemented by the map $E_{\rho_{0},\mathrm{S}}$. Infinitely many other choices are possible and become inequivalent whenever $\bar{\rho}_{i}$ and $\rho_{0}$ do not commute, leading to different lower bounds; see App. C. This intuitive reasoning was originally identified by Personick [personick1971application] and is semantically consistent with local QET, in which the logarithmic derivative of a probability distribution admits inequivalent operator generalisations in the quantum setting [Hayashi2017c].
Crucially, the SPM bound has the additional property of being attainable in single-parameter estimation by projecting onto the eigenbasis of the single SPM operator available in that case [personick1971application, rubio2020bayesian, demkowicz2020multi, rubio2023quantum]. Consequently, upon performing the eigendecomposition
@@
for the $i$-th operator, where we assume a discrete eigenbasis for simplicity but without loss of generality, the spectrum $\{s_{i,n}\}$ defines the estimator (see Refs. [personick1971application, rubio2023quantum] for details, including a continuous-variable formulation)
@@
While the PM estimators remain optimal in the multiparameter case [rubio2020bayesian], the SPM bound is generally not attainable due to the non-commutativity of the SPM operators. Nevertheless, it provides a valid lower bound on the minimum MSL and can be interpreted as the optimum obtained when incompatibility among the individually optimal measurements is neglected. Accordingly, the next section introduces a notion of measurement incompatibility based on $\mathcal{L}_{\mathrm{SPM}}$, establishing it as a result of Bayesian estimation theory analogous in fundamental significance to the Helstrom (SLD) CRB in local QET [demkowicz2020multi].
@@ III.1 Incompatibility figure of merit
where the inequality follows from the SPM lower bound in Eq. (21). A similar figure of merit has been studied also in the framework of local QET [Belliardo2021]. In the following sections, we shall derive upper and lower bounds on $\mathcal{I}$ by obtaining corresponding bounds on the unknown $\mathcal{L}_{\mathrm{min}}$.
@@ III.2.1 Pretty good measurement
Taking inspiration from hypothesis testing, we introduce the square root measurement, also known as the pretty good measurement (PGM) [Belavkin1975a, Belavkin1975, Kholevo1979, Hausladen1994]:
@@
This measurement is usually introduced for discriminating among discrete hypotheses, where the Barnum–Knill theorem guarantees the “pretty good” performance [Barnum2002]. Recently, this theorem has been extended to continuous hypotheses [Mishra2025] and applied to prove that the MSL of the PGM is within a factor two of the optimal $\mathcal{L}_{\mathrm{min}}$. While this remarkable result establishes the usefulness of the PGM for parameter estimation, it does not shed light on the issue of measurement incompatibility, i.e., understanding the discrepancy between $\mathcal{L}_{\mathrm{min}}$ and $\mathcal{L}_{\mathrm{SPM}}$. This will be the precise topic of the next subsection. In the remainder of this subsection, we set the stage by deriving the expression for the MSL achieved by the PGM.
Before proceeding, it is important to emphasise that the PGM is a POVM with outcomes in the parameter space and therefore already incorporates a classical estimator, which may not be optimal. Nonetheless, it yields simple expressions for the MSL that are easy to interpret and compare with the baseline provided by the SPM bound. In Sec. III.2.3 we will show how the the PM estimator—the optimal choice—can achieve a better MSL, at the expense of having a more involved expression.
@@
This map, like $E_{\rho,S}$ in Eq. (23), is one of the infinitely many inequivalent ways of introducing a superoperator representing “multiplication by $\rho$”. This plurality is due to noncommutativity in quantum theory, and indeed for diagonal $\rho$ and $X$, i.e., embedding classical probability into quantum objects, they all correspond to an element-wise multiplication of probability vectors. Such maps are fundamental mathematical objects in the theory of monotone metrics [Petz1996a, Hayashi2017c, Scandi2025].
@@ III.2.2 Fundamental incompatibility upper bound
We remark that the upper bound in Eq. (35) is our second and one of our key results. Interestingly, this mirrors a well-known result in local QET: the HCRB is at most twice the Helstrom (SLD) CRB [Tsang2019, Carollo2019]. Notice, however, that the HCRB is in general attainable only asymptotically with collective measurements on multiple copies. Hence, only the fundamental asymptotic impact of measurement incompatibility is limited by the upper bound on the HCRB [Belliardo2021, Das2024]. On the other hand, our result is obtained in a Bayesian estimation setting, and holds for a single shot.
@@ III.2.3 Pretty good measurement with optimal estimator
Such a behaviour may appear counter-intuitive, since measuring a system is supposed to increase the information about the parameters’ values. The intuition is re-established if one realises that this happens only because of a suboptimal choice of the classical estimator in the construction of the PGM bound. Similar situations may appear also in quantum state discrimination, i.e., for discrete hypotheses. Indeed, it has even been shown that, for specific ensembles, guessing the hypothesis that maximises the prior probability is optimal and the probability of success cannot be improved upon by performing measurements on the system [Hunter2003, Barnett2009a].
@@
Contrary to the standard PGM, this quantity is guaranteed to be smaller than $\mathcal{L}_{\mathrm{prior}}$, which can be seen explicitly by applying Jensen’s inequality to the second term of Eq. (39), since the function of $\rho(\boldsymbol{\theta})$ is quadratic and convex. Whilst for an arbitrary POVM with the same estimator—which is guaranteed to be smaller than $\mathcal{L}_{\mathrm{prior}}$—the MSL may be larger than $\mathcal{L}_{\mathrm{PGM}}$. Note that, as discussed, for our choice of quadratic loss we have a formal guarantee that the PM is the optimal classical estimator, but there is numerical evidence that it also leads to a better fidelity than the standard PGM approach in state estimation [Quadeer2024].
@@ III.3 Lower bounds
Having established an upper bound on the MSL to characterise the worst-case impact of measurement incompatibility, as well as a tighter version, our aim is now to tighten the lower bound in Eq. (36) for a given model. To this end, we seek lower bounds on the MSL that are tighter than $\mathcal{L}_{\mathrm{SPM}}$ in the presence of non-commuting SPM operators. In this section, we focus primarily on the tightest known lower bound on the MSL available in the literature, the Nagaoka–Hayashi (NH) bound, $\mathcal{L}_{\mathrm{NH}}$, recently introduced in Bayesian QET by Suzuki [Suzuki2024a].
@@ III.3.1 Nagaoka–Hayashi bound
To derive this bound, one starts by noticing that any choice of POVM and estimators satisfies the following inequality [Suzuki2024a, Lemma 1],
@@
The bound $\mathcal{L}_{\mathrm{NH}}$ formally mirrors the NH Cramér-Rao bound (NHCRB) in local QET [Conlon2020], and analogously it can be evaluated by solving a semidefinite program (SDP) [Suzuki2024a]. Notice, however, that the upper bound in Eq. (35) that holds in the Bayesian case has no direct counterpart in the local case, since the NHCRB can be much larger than twice the Helstrom (SLD) CRB, and this gap can grow with the dimension of the Hilbert space and number of parameters [Das2024].
For two parameters, the optimisation over the second moment operators $M_{11}$, $M_{12}$ and $M_{22}$ can be solved exactly. For the sake of clarity, but without loss of generality [^4] Similarly to local QET, the scalar NH bound can be recast to an equivalent problem with $L=\mathbb{I}$ and suitably redefined first moment operators $\bar{\rho}_{i}^{\prime}=\sum_{j}(\sqrt{L})^{ij}\bar{\rho}_{j}$. The explicit expression for a general weight matrix is also reported in Ref. [Suzuki2024a, Eq. (20)]., we present this expression for an identity weight matrix:
@@
where $\|A\|_{1}=\operatorname{Tr}[\sqrt{A^{\dagger}A}]$ denotes the trace norm, i.e., the sum of the singular values. This expression can be derived from Eq. (43), following, e.g., Ref. [Conlon2020, Supplementary Note 2]. This form is the one originally obtained by Nagaoka for local two-parameter estimation [Nagaoka1989].
@@ III.3.2 Other lower bounds
We first mention the Holevo bound, here denoted as $\mathcal{L}_{\mathrm{H}}$, which is a Bayesian version of the HCRB [Suzuki2024a, Zhang2024g]. It can be obtained by loosening the constraints in Eq. (42) taking the expectation value with respect to $\rho_{0}$, thereby obtaining a minimisation problem with an inequality constraint involving $d{\times}d$ matrices. The HCRB sits strictly between the SPM and NH bounds: $\mathcal{L}_{\mathrm{NH}}\geq\mathcal{L}_{\mathrm{H}}\geq\mathcal{L}_{\mathrm{SPM}}$. While in local QET the HCRB has a fundamental meaning, being asymptotically attainable, the Bayesian version does not seem to share the same significance. [^5] Note that such a Bayesian Holevo-type bound is different from a so-called Bayesian CRBs obtained by averaging the HCRB over the prior [Gill2008], conjectured to be attainable in the large sample limit of Bayesian estimation [demkowicz2020multi, Sec. 5.5], see the discussion in Sec. V. The Bayesian Holevo bound may be computationally more efficient because, while it can be computed by solving an SDP as the NH bound, it involves positive semidefinite constraints in terms of matrices of smaller dimension [Suzuki2024a].
Pursuing further the goal of identifying easily computable bounds, one may instead seek lower bounds that capture measurement incompatibility to some extent and can always be expressed in closed form, rather than indirectly as the outcome of an optimisation. A whole class of such bounds exists, based on the idea of non-commutative PM operators, of which SPMs are just one example. This is in full analogy with logarithmic derivative operators in local QET [Hayashi2017c, Ch. 6]. In App. C, we derive this family of lower bounds in a unified manner, employing the theory of monotone metrics, and generalising the approach of Ref. [rubio2020bayesian]. Here, we only mention the right posterior mean (RPM) bound
@@
written in terms of the complex matrix $\mathcal{K}_{\mathrm{R}}$. Such a matrix is analogous to the right logarithmic derivative quantum Fisher information matrix in local QET. This bound is less informative than the NH or Holevo ones, since $\mathcal{L}_{\mathrm{H}}\geq\mathcal{L}_{\mathrm{RPM}}$ [Zhang2024g, Thm. 8]. However, it can be tighter or looser than $\mathcal{L}_{\mathrm{SPM}}$ depending on the specific model. For example, it is always less tight than the SPM bound for $d=1$, where the latter is attainable. Conversely, Holevo has shown the RPM bound to be attainable for a multiparameter problem with non-commuting SPM operators: estimation of the average position and momentum of a Gaussian state with a Gaussian prior [kholevo1978investigations].

### v2

@@ II.3 Optimal quantum strategies
which differs from the a priori MSL for nonvanishing prior mean, i.e., when $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,\theta_{i}\neq 0$ for at least some $i$; see Eq. (37) and App. A. The element $L^{ij}$ denote the $ij$-th component of $L$, and we employ the Einstein summation convention over indices $i,j=1,\dots,d$.
@@
where to obtain the rightmost expressions we have combined the estimators and the measurement by using the relabelled POVM elements [64]
@@ II.4 Lower bounds from optimal single-parameter strategies
While exact computation of $\mathcal{L}_{\mathrm{min}}$ is generally difficult, it admits a lower bound [40, 126, 64]
@@
Here and in the rest of this work, we assume that this inverse exists, which is guaranteed for a full-rank $\rho_{0}$. When this is not the case, one simply needs to work in the relevant subspace spanned by the parametric family $\rho(\boldsymbol{\theta})$, which corresponds to the support of $\rho_{0}$. [^3] This is possible because the support the operators $\{\bar{\rho}_{j}\}$ is contained in the support of $\rho_{0}$, since $\braket{\psi|\rho_{0}|\psi}=0$ implies $\braket{\psi|\rho(\boldsymbol{\theta})|\psi}=0$ almost everywhere (in the measure-theoretic sense), which in turn implies $\braket{\psi|\bar{\rho}_{j}|\psi}=0$. This is markedly different from local QET, where the partial derivatives operators may have components also in the kernel of the state, and properties of rank-deficient models are non-trivial. For example, necessary and sufficient conditions for attainability of the SLD CRB are not fully characterized [127, 128, 89, 129] (except for pure states [78, 130])
@@
Formally, the PM estimator is the component-wise ratio of the first-moment probability vector $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,p(x|\boldsymbol{\theta})\,\theta_{i}$ to the average probability vector $\int\mathrm{d}\boldsymbol{\theta}\,p(\boldsymbol{\theta})\,p(x|\boldsymbol{\theta})$. Intuitively, if $p(x|\boldsymbol{\theta})$ is formally replaced by a density operator $\rho(\boldsymbol{\theta})$, there are infinitely many inequivalent ways to “divide” $\bar{\rho}_{i}$ by $\rho_{0}$. The SPMs correspond to the symmetric choice, where “multiplication by $\rho_{0}$” is implemented by the map $E_{\rho_{0},\mathrm{S}}$. Infinitely many other choices are possible and become inequivalent whenever $\bar{\rho}_{i}$ and $\rho_{0}$ do not commute, leading to different lower bounds; see App. C. This intuitive reasoning was originally identified by Personick [23] and is semantically consistent with local QET, in which the logarithmic derivative of a probability distribution admits inequivalent operator generalisations in the quantum setting [110].
Crucially, the SPM bound has the additional property of being attainable in single-parameter estimation by projecting onto the eigenbasis of the single SPM operator available in that case [23, 40, 64, 131]. Consequently, upon performing the eigendecomposition
@@
for the $i$-th operator, where we assume a discrete eigenbasis for simplicity but without loss of generality, the spectrum $\{s_{i,n}\}$ defines the estimator (see Refs. [23, 131] for details, including a continuous-variable formulation)
@@
While the PM estimators remain optimal in the multiparameter case [40], the SPM bound is generally not attainable due to the non-commutativity of the SPM operators. Nevertheless, it provides a valid lower bound on the minimum MSL and can be interpreted as the optimum obtained when incompatibility among the individually optimal measurements is neglected. Accordingly, the next section introduces a notion of measurement incompatibility based on $\mathcal{L}_{\mathrm{SPM}}$, establishing it as a result of Bayesian estimation theory analogous in fundamental significance to the Helstrom (SLD) CRB in local QET [64].
@@ III.1 Incompatibility figure of merit
where the inequality follows from the SPM lower bound in Eq. (21). A similar figure of merit has been studied also in the framework of local QET [61]. In the following sections, we shall derive upper and lower bounds on $\mathcal{I}$ by obtaining corresponding bounds on the unknown $\mathcal{L}_{\mathrm{min}}$.
@@ III.2.1 Pretty good measurement
Taking inspiration from hypothesis testing, we introduce the square root measurement, also known as the pretty good measurement (PGM) [106, 107, 108, 109]:
@@
This measurement is usually introduced for discriminating among discrete hypotheses, where the Barnum–Knill theorem guarantees the “pretty good” performance [132]. Recently, this theorem has been extended to continuous hypotheses [133] and applied to prove that the MSL of the PGM is within a factor two of the optimal $\mathcal{L}_{\mathrm{min}}$. While this remarkable result establishes the usefulness of the PGM for parameter estimation, it does not shed light on the issue of measurement incompatibility, i.e., understanding the discrepancy between $\mathcal{L}_{\mathrm{min}}$ and $\mathcal{L}_{\mathrm{SPM}}$. This will be the precise topic of the next subsection. In the remainder of this subsection, we set the stage by deriving the expression for the MSL achieved by the PGM.
Before proceeding, it is important to emphasise that the PGM is a POVM with outcomes in the parameter space and therefore already incorporates a classical estimator, which may not be optimal. Nonetheless, it yields simple expressions for the MSL that are easy to interpret and compare with the baseline provided by the SPM bound. In Sec. III.2.3 we will show how the PM estimator—the optimal choice—can achieve a better MSL, at the expense of having a more involved expression.
@@
This map, like $E_{\rho,S}$ in Eq. (23), is one of the infinitely many inequivalent ways of introducing a superoperator representing “multiplication by $\rho$”. This plurality is due to noncommutativity in quantum theory, and indeed for diagonal $\rho$ and $X$, i.e., embedding classical probability into quantum objects, they all correspond to an element-wise multiplication of probability vectors. Such maps are fundamental mathematical objects in the theory of monotone metrics [134, 110, 135].
@@ III.2.2 Fundamental incompatibility upper bound
We remark that the upper bound in Eq. (35) is our second and one of our key results. Interestingly, this mirrors a well-known result in local QET: the HCRB is at most twice the Helstrom (SLD) CRB [76, 77]. Notice, however, that the HCRB is in general attainable only asymptotically with collective measurements on multiple copies. Hence, only the fundamental asymptotic impact of measurement incompatibility is limited by the upper bound on the HCRB [61, 91]. On the other hand, our result is obtained in a Bayesian estimation setting, and holds for a single shot.
@@ III.2.3 Pretty good measurement with optimal estimator
Such a behaviour may appear counter-intuitive, since measuring a system is supposed to increase the information about the parameters’ values. The intuition is re-established if one realises that this happens only because of a suboptimal choice of the classical estimator in the construction of the PGM bound. Similar situations may appear also in quantum state discrimination, i.e., for discrete hypotheses. Indeed, it has even been shown that, for specific ensembles, guessing the hypothesis that maximises the prior probability is optimal and the probability of success cannot be improved upon by performing measurements on the system [136, 137].
@@
Contrary to the standard PGM, this quantity is guaranteed to be smaller than $\mathcal{L}_{\mathrm{prior}}$, which can be seen explicitly by applying Jensen’s inequality to the second term of Eq. (39), since the function of $\rho(\boldsymbol{\theta})$ is quadratic and convex. Whilst for an arbitrary POVM with the same estimator—which is guaranteed to be smaller than $\mathcal{L}_{\mathrm{prior}}$—the MSL may be larger than $\mathcal{L}_{\mathrm{PGM}}$. Note that, as discussed, for our choice of quadratic loss we have a formal guarantee that the PM is the optimal classical estimator, but there is numerical evidence that it also leads to a better fidelity than the standard PGM approach in state estimation [138].
@@ III.3 Lower bounds
Having established an upper bound on the MSL to characterise the worst-case impact of measurement incompatibility, as well as a tighter version, our aim is now to tighten the lower bound in Eq. (36) for a given model. To this end, we seek lower bounds on the MSL that are tighter than $\mathcal{L}_{\mathrm{SPM}}$ in the presence of non-commuting SPM operators. In this section, we focus primarily on the tightest known lower bound on the MSL available in the literature, the Nagaoka–Hayashi (NH) bound, $\mathcal{L}_{\mathrm{NH}}$, recently introduced in Bayesian QET by Suzuki [102].
@@ III.3.1 Nagaoka–Hayashi bound
To derive this bound, one starts by noticing that any choice of POVM and estimators satisfies the following inequality [102, Lemma 1],
@@
The bound $\mathcal{L}_{\mathrm{NH}}$ formally mirrors the NH Cramér-Rao bound (NHCRB) in local QET [88], and analogously it can be evaluated by solving a semidefinite program (SDP) [102]. Notice, however, that the upper bound in Eq. (35) that holds in the Bayesian case has no direct counterpart in the local case, since the NHCRB can be much larger than twice the Helstrom (SLD) CRB, and this gap can grow with the dimension of the Hilbert space and number of parameters [91].
For two parameters, the optimisation over the second moment operators $M_{11}$, $M_{12}$ and $M_{22}$ can be solved exactly. For the sake of clarity, but without loss of generality [^4] Similarly to local QET, the scalar NH bound can be recast to an equivalent problem with $L=\mathbb{I}$ and suitably redefined first moment operators $\bar{\rho}_{i}^{\prime}=\sum_{j}(\sqrt{L})^{ij}\bar{\rho}_{j}$. The explicit expression for a general weight matrix is also reported in Ref. [102, Eq. (20)]., we present this expression for an identity weight matrix:
@@
where $\|A\|_{1}=\operatorname{Tr}[\sqrt{A^{\dagger}A}]$ denotes the trace norm, i.e., the sum of the singular values. This expression can be derived from Eq. (43), following, e.g., Ref. [88, Supplementary Note 2]. This form is the one originally obtained by Nagaoka for local two-parameter estimation [85].
@@ III.3.2 Other lower bounds
We first mention the Holevo bound, here denoted as $\mathcal{L}_{\mathrm{H}}$, which is a Bayesian version of the HCRB [102, 103]. It can be obtained by loosening the constraints in Eq. (42) taking the expectation value with respect to $\rho_{0}$, thereby obtaining a minimisation problem with an inequality constraint involving $d{\times}d$ matrices. The HCRB sits strictly between the SPM and NH bounds: $\mathcal{L}_{\mathrm{NH}}\geq\mathcal{L}_{\mathrm{H}}\geq\mathcal{L}_{\mathrm{SPM}}$. While in local QET the HCRB has a fundamental meaning, being asymptotically attainable, the Bayesian version does not seem to share the same significance. [^5] Note that such a Bayesian Holevo-type bound is different from a so-called Bayesian CRBs obtained by averaging the HCRB over the prior [139], conjectured to be attainable in the large sample limit of Bayesian estimation [64, Sec. 5.5], see the discussion in Sec. V. The Bayesian Holevo bound may be computationally more efficient because, while it can be computed by solving an SDP as the NH bound, it involves positive semidefinite constraints in terms of matrices of smaller dimension [102].
Pursuing further the goal of identifying easily computable bounds, one may instead seek lower bounds that capture measurement incompatibility to some extent and can always be expressed in closed form, rather than indirectly as the outcome of an optimisation. A whole class of such bounds exists, based on the idea of non-commutative PM operators, of which SPMs are just one example. This is in full analogy with logarithmic derivative operators in local QET [110, Ch. 6]. In App. C, we derive this family of lower bounds in a unified manner, employing the theory of monotone metrics, and generalising the approach of Ref. [40]. Here, we only mention the right posterior mean (RPM) bound
@@
written in terms of the complex matrix $\mathcal{K}_{\mathrm{R}}$. Such a matrix is analogous to the right logarithmic derivative quantum Fisher information matrix in local QET. This bound is less informative than the NH or Holevo ones, since $\mathcal{L}_{\mathrm{H}}\geq\mathcal{L}_{\mathrm{RPM}}$ [103, Thm. 8]. However, it can be tighter or looser than $\mathcal{L}_{\mathrm{SPM}}$ depending on the specific model. For example, it is always less tight than the SPM bound for $d=1$, where the latter is attainable. Conversely, Holevo has shown the RPM bound to be attainable for a multiparameter problem with non-commuting SPM operators: estimation of the average position and momentum of a Gaussian state with a Gaussian prior [125].